# Online Makespan Scheduling under Scenarios

### v1

@@ 3 Near-Tight Bounds for $m=2$ Machines
The proxy competitive ratio is an upper bound on the competitive ratio. The lower bounds that we use are already well-known; e.g., Graham’s List Scheduling Algorithm is analyzed by bounding the proxy competitive ratio from above. We take this approach, however, one step further and restrict ourselves to analyzing only some instances that attain the largest possible proxy competitive ratio (see next Observation). This is significantly more difficult for the competitive ratio in the usual sense, as a witness to having good or bad offline optima highly depends on how the processing times of jobs fit together numerically.
Observation 3.4 (cf. Figure 1).
@@
Let $\mathcal{A}$ be an online algorithm for OMSS(2,2)22(2,2)( 2 , 2 ) obeying Rule 3.3. Assume that $\mathcal{A}$ has a proxy competitive ratio of at most $\frac{5}{3}$ on instances where
@@
Then, $\mathcal{A}$ is $\frac{5}{3}$-competitive.
The above observation motivates the following parameter of significant value.
@@
Figure 1: Illustrative description of Observation 3.4. Processing times $p_{j}$ are given by the widths. Given the dark blue schedule, a worst outcome turns out to be the light blue job $n-1$ with a processing time $p_{n-1}$ such that $C_{n-1}=C_{j}$, followed by the light red job $n$ with $p_{n}=p(S_{1}\cap[n-1])$.
@@
$\alpha\coloneqq\begin{cases}\frac{p(J_{1}\cap S_{1})}{\max\left\{p(J_{1}\cap S_{2}),p(J_{2}\cap S_{1})+p(J_{1}\cap S_{1})-p(J_{2}\cap S_{2})\right\}}&\text{if }p(J_{2}\cap S_{2})>p(J_{2}\cap S_{1}),\\ 0&\text{ else.}\end{cases}$
@@
In Figure 1, we have $p(J_{2}\cap S_{2}\cap[n-2])>p(J_{2}\cap S_{1}\cap[n-2])$ and the second term of the denominator corresponds to the load $p(J_{2}\cap S_{1}\cap[n-1])$, i.e., the load of $J_{2}\cap S_{1}$ after the $(n-1)$-st job is assigned. In general, anticipation measures the minimum discrepancy of a machine with respect to different scenarios after machine $2$ is loaded with double-scenario jobs just enough to attain the makespan of the schedule together with machine $1$. In our algorithm, one of the goals is to keep $\alpha\leq 2$ in order to brace ourselves for upcoming double-scenario jobs. The invariant which we maintain to this end is the following.
Invariant 3.5.
@@
Algorithm 1 shows how a single job $j$ is assigned given a schedule of the first $j-1$ jobs where Invariant 3.5 is maintained. To analyze the algorithm, we prove that instances satisfying the properties in Observation 3.4 are always assigned so that the first $n-1$ jobs maintain Invariant 3.5 and the last job results in a proxy competitive ratio of at most $\nicefrac{{5}}{{3}}$.
Algorithm 1 The assignment of a single job $j$ given a $\frac{5}{3}$-competitive schedule that satisfies Invariant 3.5.
@@
6:     if assigning $j$ to the first machine satisfies Invariant 3.5 then

### v2

@@ 3.1 Preparation for the Algorithm
The proxy competitive ratio is an upper bound on the competitive ratio. The lower bounds that we use are already well-known; for instance, Graham’s List Scheduling Algorithm is in fact analyzed by bounding the proxy competitive ratio from above. We take this approach, however, one step further and restrict ourselves to analyzing only the instances that attain the largest possible proxy competitive ratio. This would have been significantly more difficult for the competitive ratio in the usual sense, as a witness to having good or bad offline optima highly depends on how the processing times of jobs fit together numerically.
Transforming to a worst-case instance.
In search of instances with the worst possible proxy competitive ratio, we transform arbitrary instances to those that admit a certain structure, and are no better in terms of competitiveness, through a few simple operations.
Definition 3.
Let $\mathcal{I}=((n,(p_{j})_{j\in[n]},\{S_{k}\}_{k\in[K]}))$ be an instance of OMSS(2,2)(2,2). The deletion of a job $j$ is the sole modification that $p_{j}$ is set to 0, i.e., it gives rise to an instance $\mathcal{I}^{\prime}=(n,(p^{\prime}_{j})_{j\in[n]},\{S_{k}\}_{k\in[K]})$ with
$p^{\prime}_{j^{\prime}}=\begin{cases}0,&j=j^{\prime}\\ p_{j},&j\neq j^{\prime}.\end{cases}$
Lemma 3.4 (Deletion lemma, cf. Figure 1).
Let $\mathcal{I}$ be an instance on which an algorithm $\mathcal{A}$ admits a competitive ratio $\rho$. Let $j\in[n]$ be the largest index such that $j\in S_{1}\triangle S_{2}$ holds. Assume furthermore that the assignment of the $j$-th job to the $i=\tau(j)$-th machine was executed so that the scenario more represented in the machine did not change after the $j$-th job, i.e.,
$\mathrm{argmax}_{k\in\{1,2\}}p(J_{i}\cap S_{k}\cap[j-1])=\mathrm{argmax}_{k\in\{1,2\}}p(J_{i}\cap S_{k}\cap[j]).$
Then the instance $\mathcal{I}^{\prime}$ obtained by the deletion of $j$ admits a (proxy) competitive ratio $\rho^{\prime}\geq\rho.$
Proof 3.5.
Let $\tau,\tau^{\prime}\colon[n]\to\{1,2\}$ denote the assignment of the original instance and and the modified instance, respectively. Both instances admit the identical assignment up to (and including) job j1j-1. We observe that
$\mathrm{argmax}_{k\in\{1,2\}}p(J_{i}\cap S_{k}\cap[j-1])=\mathrm{argmax}_{k\in\{1,2\}}p(J_{i}\cap S_{k})$
for $i\in\{1,2\}$ for $\mathcal{I}$ and analogously for $\mathcal{I}^{\prime}$, since $\{j+1,\ldots,n\}\subseteq S_{1}\cap S_{2}$ by the choice of $j$ and therefore the difference $p(J_{i}\cap S_{1})-p(J_{1}\cap S_{2})$ stays constant after the assignment of the $j$-th job. For both instances, the jobs $j+1,\ldots,n$ are distributed to a machine $i$ such that the other machine 3i3-i admits makespan by the same reasoning. Since the assignment of the job $j$ does not affect $\max_{k\in\{1,2\}}p(J_{i}\cap S_{k}\cap[j^{\prime}])$ for any $j^{\prime}\in[n]$, we have $\tau(j^{\prime})=\tau^{\prime}(j^{\prime})$ for $j^{\prime}>j$.
The $j$-th job in instance $\mathcal{I}$ cannot be in a combination of machine and scenario that admits the makespan of the schedule. Thus the makespan $\operatorname{\mathrm{onl}}_{\mathcal{I}}$ and $\operatorname{\mathrm{onl}}_{\mathcal{I}^{\prime}}$ of both instances coincide. On the other hand, the offline optimal value or the denominator of the proxy competitive ratio do not increase by the modification, so that $\rho^{\prime}\geq\rho$ holds.
[…]
Figure 1: Deletion of the largest $j\in S_{1}\triangle S_{2}$ (checkered) does not increase the approximation ratio, as it does not affect the execution of Rule 3.3.
We may assume for the completion time of the largest job index $j\in S_{1}\triangle S_{2}$ that $C_{j}>C_{j-1}$, for otherwise, we may apply Lemma 3.4 successively to eliminate jobs $j\in S_{1}\triangle S_{2}$ that cause a violation of this inequality and analyze the resulting instances instead.
Next, we observe that under certain circumstances, we may “cut” the double-jobs $j$ in that we replace them by other jobs $j^{\prime},j^{\prime\prime}\in S_{1}\cap S_{2}$ such that $p_{j^{\prime}}+p_{j^{\prime\prime}}=p_{j}$, while ensuring that the behavior of our algorithm remains the same way:
Lemma 3.6 (Cutting lemma, cf. Figure 2).
Let $t\in S_{1}\cap S_{2}$ be a job index for an instance $\mathcal{I}=(n,(p_{j})_{j\in[n]},\{S^{\prime}_{k}\}_{k\in\{1,2\}})$ of OMSS(2,2) such that $\mathrm{makespan}(t)>\mathrm{makespan}(t-1)$. Consider a schedule $\tau\colon[n]\to\{1,2\}$ of $n$ jobs, on which the algorithm admits a (proxy) competitive ratio of $\rho$. Then there is another instance $\mathcal{I}^{\prime}=(n+1,(p^{\prime}_{j})_{j\in[n]},\{S^{\prime}_{k}\}_{k\in\{1,2\}})$ with $n+1$ jobs with processing times $p^{\prime}_{j},j\in[n+1]$ under scenarios $S_{1}^{\prime},S_{2}^{\prime}\subseteq[n+1]$ with the following properties:
  1. (i)
    $p^{\prime}(J^{\prime}_{i}\cap S^{\prime}_{k}\cap[j])=p(J_{i}\cap S_{k}\cap[j])$ for $j\in[t-1]$, $i\in\{1,2\}$ and $k\in\{1,2\}$.
  2. (ii)
    The algorithm assigns the first $t$ jobs of $\mathcal{I}^{\prime}$ so that
    $\max_{k\in\{1,2\}}p(J^{\prime}_{1}\cap S^{\prime}_{k}\cap[t])=\max_{k\in\{1,2\}}p(J^{\prime}_{2}\cap S^{\prime}_{k}\cap[t]).$
  3. (iii)
    The algorithm admits a (proxy) competitive ratio $\rho^{\prime}\geq\rho$ on the instance $\mathcal{I}^{\prime}$.
Proof 3.7.
We define the new instance $\mathcal{I}^{\prime}$ as follows: we set
$S^{\prime}_{k}=(S_{k}\cap[t])\;\dot{\cup}\;[t+1]\;\dot{\cup}\;\{j\in\{t+2,\ldots,n+1\}\colon j-1\in S_{k}\}$
for $t\in[n]$ and $k\in\{1,2\}$. Further, we define $p^{\prime}_{j}=p_{j}$ for $1\leq j<t$ and $p^{\prime}_{j}=p_{j-1}$ for $t+2\leq j\leq n+1$. We determine $p^{\prime}_{t}$ and $p^{\prime}_{t+1}$ later. We will aim to satisfy $p^{\prime}_{t}+p^{\prime}_{t+1}=p_{t}$.
We start by noticing that the instances $\mathcal{I}$ and $\mathcal{I}^{\prime}$ coincide when restricted to the first t1t-1 jobs. Therefore, the algorithm will yield identical assignments until the assignment of the $t$-th job and the first property is satisfied.
Now we consider the job indexed $t$. In both instances, we have $t\in S_{1}\cap S_{2}$ resp. $t\in S_{1}^{\prime}\cap S_{2}^{\prime}$. For such a double-scenario job, the assignment only depends on the prior subschedule $\tau|_{[t-1]}=\tau^{\prime}|_{[t-1]}$ and not on the processing time $p_{t}$ resp. $p_{t}^{\prime}$. Therefore, we also have $\tau(t)=\tau(t^{\prime})$ for any $p^{\prime}_{t}\in(0,p_{t})$. To obtain the second property, we pick
$p^{\prime}_{t}=\mathrm{makespan}(t-1)-\min_{i^{\prime}\in\{1,2\}}\max_{k^{\prime}\in\{1,2\}}p(J^{\prime}_{i^{\prime}}\cap S^{\prime}_{k}\cap[t-1]),$ (1)
and $p^{\prime}_{t+1}=p_{t}-p^{\prime}_{t}$. We guarantee that $p^{\prime}_{t+1}>0$ because of $\mathrm{makespan}(t)>\mathrm{makespan}(t-1)$. Since the $t$-th job is assigned to the $i^{\prime}$-th machine attaining the minimum in (1), we then have
$\max_{k\in\{1,2\}}p^{\prime}(J^{\prime}_{i^{\prime}}\cap S^{\prime}_{k}\cap[t])=\max_{k\in\{1,2\}}p^{\prime}(J^{\prime}_{i^{\prime}}\cap S^{\prime}_{k}\cap[t-1])+p^{\prime}_{t}=\mathrm{makespan}(t-1).$ (2)
Since the $t$-th job was not assigned to the $(3-i^{\prime})$-th machine, we also have
$\max_{k\in\{1,2\}}\{p^{\prime}(J^{\prime}_{3-i^{\prime}}\cap S^{\prime}_{k}\cap[t])\}=\max_{k\in\{1,2\}}\{p^{\prime}(J^{\prime}_{3-i^{\prime}}\cap S^{\prime}_{k}\cap[t-1])\}=\mathrm{makespan}(t-1).$ (3)
Comparing lines (2) and (3) reveals that the second property is satisfied.
For the third property, we observe that both machines admit makespan prior to the insertion of the $(t+1)$-st job, so the algorithm will assign this job to the machine where it has assigned the previous job $t$ to. This yields
$p^{\prime}(J^{\prime}_{i^{\prime}}\cap S^{\prime}_{k}\cap[t+1])$ $=p^{\prime}(J^{\prime}_{i^{\prime}}\cap S^{\prime}_{k}\cap[t-1])+p^{\prime}_{t}+p^{\prime}_{t+1}$
$=p(J_{i^{\prime}}\cap S_{k}\cap[t-1])+p_{t}=p(J_{i^{\prime}}\cap S_{k}\cap[t])$
for $k\in\{1,2\}$. Since the $j$-th job in $\mathcal{I}$ is identical to the $(j+1)$-st job in $\mathcal{I}^{\prime}$, the algorithm will output the same value in both cases. It is also straightforward to observe that the denominator of the proxy competitive ratio cannot have increased after the modification.
In total, this implies $\rho^{\prime}\geq\rho,$ proving that the third property is satisfied.
[…]
Figure 2: Execution of the cutting lemma on job $n$, before (left) and after (right).
We observe that in order to analyze the proxy competitive ratio of instances whose last job $n$ satisfies $n\in S_{1}\cap S_{2}$, it suffices to consider those instances, in which additionally the first n1n-1 jobs satisfy
$p(S_{k}\cap J_{k}\cap[n-1])=p(S_{3-k}\cap J_{3-k}\cap[n-1]).$ (4)
Indeed, if we have an arbitrary instance with last job $n\in S_{1}\cap S_{2}$, we can apply Lemma 3.6 with $t=n$.
On the other hand, if (4) is satisfied, the adversarial last job $n$ is double-scenario and has processing time
$p_{n}=p(S_{k}\cap J_{k}\cap[n-1])+\max_{k\in\{1,2\}}p(S_{k}\cap J_{3-k}\cap[n-1])=\max_{k\in\{1,2\}}p(S_{k}\cap[n-1])$
, so as to maximize the proxy competitive ratio
$\rho$ $=\frac{p(S_{k}\cap J_{k}\cap[n-1])+p_{n}}{\max\left\{\max_{k\in\{1,2\}}\left\{\frac{p(S_{k}\cap[n-1])+p_{n}}{2}\right\},p_{n}\right\}}$ (5)
$=\frac{2p(S_{k}\cap J_{k}\cap[n-1])+\max_{k^{\prime}\in\{1,2\}}p(S_{k^{\prime}}\cap J_{3-k^{\prime}}\cap[n-1])}{p(S_{k}\cap J_{k}\cap[n-1])+\max_{k^{\prime}\in\{1,2\}}p(S_{k^{\prime}}\cap J_{3-k^{\prime}}\cap[n-1])}$ (6)
$=1+\frac{p(S_{k}\cap J_{k}\cap[n-1])}{\max_{k^{\prime}\in\{1,2\}}p(S_{k^{\prime}}\cap J_{3-k^{\prime}}\cap[n-1])}.$ (7)
Now let $j\in[n]$ be the highest-indexed job that is in $S_{1}\triangle S_{2}$. If $j=n$, we assume to be done, so let $j<n$. By Lemma 3.4, there is a job $j^{\prime}>j$ whose completion time is larger than that of $j$ in some scenario. We can pick the smallest-indexed $j^{\prime}$ whose completion time in some scenario $S_{k}$ surpasses the completion time of $j$. For this index $j^{\prime}$, we apply Lemma 3.6 so that after the assignment of the earlier of the two jobs that replace $j^{\prime}$, the completion time of the job $j$ still equals the overall makespan of the schedule. This implies:
Observation 3.8.
Among the instances $\mathcal{I}=(n,(p_{j})_{j\in[n]},\{S_{1},S_{2}\})$ with $n\in S_{1}\cap S_{2}$, there is a worst-case instance with the following properties:
  1. (i)
    If $j\in[n]$ is the largest index with $j\in S_{1}\triangle S_{2}$, there is another job indexed $j^{\prime}>j$ such that $C_{j}=C_{j^{\prime}}$.
  2. (ii)
    Equation (4) holds, i.e., $p(S_{k}\cap J_{k}\cap[n-1])=p(S_{3-k}\cap J_{3-k}\cap[n-1])$.
We transform this instance further by deleting the jobs $j^{\prime}+1,\ldots,n-1$ and changing $p_{n}$ to $p(S_{k}\cap J_{k}\cap[j^{\prime}])+\max_{k^{\prime}\in\{1,2\}}p(S_{k^{\prime}}\cap J_{3-k^{\prime}}\cap[j^{\prime}])$. Since $j^{\prime}$ is the penultimate job after the modification, the equation (4) holds for the modified instance as well.
If $\rho^{\prime}$ denotes the new proxy competitive ratio after the modifications, then we have
$\rho$ $=1+\frac{p(S_{k}\cap J_{k}\cap[n-1])}{\max_{k^{\prime}\in\{1,2\}}p(S_{k^{\prime}}\cap J_{3-k^{\prime}}\cap[n-1])}$ (8)
$=1+\frac{p(S_{k}\cap J_{k}\cap[j^{\prime}])+p(J_{k}\cap S_{k}\cap\{j^{\prime}+1,\ldots,n-1\})}{\max_{k^{\prime}\in\{1,2\}}(p(S_{k^{\prime}}\cap J_{3-k^{\prime}}\cap[j^{\prime}])+p(S_{k^{\prime}}\cap J_{3-k^{\prime}}\cap\{j^{\prime}+1\ldots,n-1\}))}$ (9)
$=1+\frac{p(S_{k}\cap J_{k}\cap[j^{\prime}])+p(J_{k}\cap\{j^{\prime}+1,\ldots,n-1\})}{\max_{k^{\prime}\in\{1,2\}}(p(S_{k^{\prime}}\cap J_{3-k^{\prime}}\cap[j^{\prime}])+p(J_{3-k^{\prime}}\cap\{j^{\prime}+1\ldots,n-1\}))}$ (10)
$\stackrel{{\scriptstyle(*)}}{{\leq}}1+\frac{p(S_{k}\cap J_{k}\cap[j^{\prime}]))}{\max_{k^{\prime}\in\{1,2\}}p(S_{k^{\prime}}\cap J_{3-k^{\prime}}\cap[j^{\prime}])}=\rho^{\prime}.$ (11)
Here, the inequality ()(*) follows by Observation 3.8. Since $\rho$ is the proxy competitive ratio of a worst-case instance among those with $n\in S_{1}\cap S_{2}$, we conclude that the modified instance is a worst-case instance as well. We can therefore further strengthen the properties sketched in Observation 3.8 by assuming that the jobs before the penultimate job $j^{\prime}=n-1$ and after the last job $j\in S_{1}\triangle S_{2}$ had smaller completion time than the job $j$. In summary, we can observe the following statement.
Observation 3.9 (cf. Figure 3).
@@
Let $\mathcal{A}$ be an online algorithm for OMSS(2,2)(2,2) obeying Rule 3.3. Assume that $\mathcal{A}$ has a proxy competitive ratio of at most $\frac{5}{3}$ on instances where
@@
Then, $\mathcal{A}$ is $\nicefrac{{5}}{{3}}$-competitive.
@@
Figure 3: Illustrative description of Observation 3.9. Given the dark blue schedule, a worst outcome turns out to be the light blue job n1n-1 with a processing time $p_{n-1}$ such that $C_{n-1}=C_{j}$, followed by the light red job $n$ with $p_{n}=p(S_{1}\cap[n-1])$. The anticipation measures the minimum discrepancy in either machine prior to the revelation of the final job $n$.
With Lemma 3.4 in mind, we can conclude: among instances with $n\in S_{1}\cap S_{2}$, at least one worst-case instance behaves as follows: The last single-scenario job $j\in S_{1}\triangle S_{2}$ admits the makespan at the time of its insertion. Then, the double-scenario jobs $j+1,\ldots,n-1$ are all assigned to the opposite machine. After the insertion of the (n1)(n-1)-st job, both machines have the same makespan; i.e., the schedule satisfies (4). The last job $n$ has the adversarial processing time of
$p_{n}=\max_{k\in\{1,2\}}p(S_{k}\cap[n-1])=p(S_{k}\cap J_{k}\cap[n-1])+\max_{k\in\{1,2\}}p(S_{k}\cap J_{3-k}\cap[n-1]).$
Figure 3 illustrates a “worst” possible outcome, given a partial schedule.
Anticipation.
In the sequel, we assume that the instance which we analyze satisfies the properties in Observation 3.9. For such a special instance, a key parameter turns out to capture discrepancies within each machine.
@@
$\alpha\coloneqq\begin{cases}\frac{p(J_{1}\cap S_{1})}{\max\left\{p(J_{1}\cap S_{2}),p(J_{2}\cap S_{1})+p(J_{1}\cap S_{1})-p(J_{2}\cap S_{2})\right\}}&\text{ if }p(J_{2}\cap S_{2})>p(J_{2}\cap S_{1})\\ 0&\text{ else.}\end{cases}$
@@
In Figure 3, we have $p(J_{2}\cap S_{2}\cap[n-2])>p(J_{2}\cap S_{1}\cap[n-2])$ and the second term of the denominator corresponds to the load $p(J_{2}\cap S_{1}\cap[n-1])$, i.e., the load of $J_{2}\cap S_{1}$ after the (n1)(n-1)-st job is assigned. To further motivate the concept of anticipation, let us bound the proxy competitive ratio in terms of it.
Lemma 3.10.
Let a schedule of n1n-1 jobs be given that satisfies (4) and anticipation $\alpha$. After any assignment of the $n$-th job, the resulting schedule has proxy competitive ratio at most $\frac{2\alpha+1}{\alpha+1}$.
Proof 3.11.
Let $P\coloneqq p(S_{1}\cap J_{1}\cap[n-1])$. Then if $p_{n}\leq P+\frac{P}{\alpha}$, we have
$\rho\leq\frac{2P+2p_{n}}{P+{\max\{p(S_{1}\cap J_{1}),p(S_{2}\cap J_{2})\}+p_{n}}}\leq\frac{2P+2p_{n}}{P+\frac{P}{\alpha}+P\left(1+\frac{1}{\alpha}\right)}\leq\frac{2\alpha+1}{\alpha+1},$
and otherwise,
$\rho\leq\frac{P+p_{n}}{p_{n}}\leq\frac{P+\left(1+\frac{1}{\alpha}\right)P}{\left(1+\frac{1}{\alpha}\right)P}=\frac{2\alpha+1}{\alpha+1},$
showing the claim.
3.2 The Algorithm
In our algorithm, one of the goals is to keep $\alpha\leq 2$ in order to brace ourselves for upcoming double-scenario jobs in order to ensure a proxy competitive ratio of at most $\nicefrac{{5}}{{3}}$. The invariant which we maintain to this end is the following.
Invariant 3.12.
@@ 3.2 The Algorithm
Algorithm 1 shows how a single job $j$ is assigned given a schedule of the first j1j-1 jobs where Invariant 3.12 is maintained. In the sequel, we prove that instances satisfying the properties in Observation 3.9 are always assigned so that the first n1n-1 jobs maintain Invariant 3.12 and the last job results in a proxy competitive ratio of at most $\nicefrac{{5}}{{3}}$.
Algorithm 1 The assignment of a single job $j$ given a $\frac{5}{3}$-competitive schedule that satisfies Invariant 3.12.
@@
6:  if assigning $j$ to the first machine satisfies Invariant 3.12 then